# In LIGO’s Sight? Vigorous Coherent Gravitational Waves from Cooled Collapsar Disks

### v1

@@ Appendix A Compact binary mergers
During the merger of a BNS or a BH-NS system, some of the NS crust is disrupted and circularized around the merger remnant. The tidal disruption of the NS naturally gives rise to non-axisymmetric modes. This phenomenon is illustrated in Figure 6(a), which presents a 3D rendering of the logarithmic mass density around the post-merger BH $10\,{\rm ms}$ after a BNS merger. The spiral structure persists for a few dozen milliseconds before transitioning into an axisymmetric configuration within a more uniformly distributed post-merger disk [Fig. 6(b)]. The compact nature of these post-merger disks (Fernández et al., 2020) indicates that their characteristic GW frequency is notably higher compared to those observed in collapsars.
Fig. 6(c-f) present the integrated GW spectra at different times after the merger. During the first $\sim 30\,{\rm ms}$, the spiral arm generates strong GW emission due to a significant deviation from axisymmetry. The contribution from different radii results in a roughly flat spectrum. As the disk starts to be accreted and become more axisymmetric, its GW emission decays with time. At $t\gtrsim 100\,\,{{\rm ms}}$, the spiral arm vanishes, and density waves at $R_{d}\approx 25\,\,{{\rm km}}$ govern the GW emission. The small radius of post-merger accretion disks results in a characteristic GW frequency of $\,{f_{\rm GW}}\approx 1.5\,{\rm kHz}$ [right panels, see Eq. (5)].

### v2

@@ Appendix A Compact binary mergers
During the merger of a BNS or a BH–NS system, some of the NS crust is disrupted and circularized around the merger remnant. The tidal disruption of the NS naturally gives rise to nonaxisymmetric modes. This phenomenon is illustrated in Figure 6(a), which presents a 3D rendering of the logarithmic mass density around the post-merger BH $10\,{\rm ms}$ after a BNS merger. The spiral structure persists for a few dozen milliseconds before transitioning into an axisymmetric configuration within a more uniformly distributed post-merger disk [Fig. 6(b)]. The compact nature of these post-merger disks (Fernández et al., 2020) indicates that their characteristic GW frequency is notably higher compared to those observed in collapsars.
Fig. 6(c-f) present the integrated GW spectra at different times after the merger. During the first $\sim 30\,{\rm ms}$, the spiral arm generates strong GW emission owing to a significant deviation from axisymmetry. The contribution from different radii results in a roughly flat spectrum. As the disk starts to be accreted and become more axisymmetric, its GW emission decays with time. At $t\gtrsim 100\,\,{{\rm ms}}$, the spiral arm vanishes, and density waves at $R_{d}\approx 25\,\,{{\rm km}}$ govern the GW emission. The small radius of post-merger accretion disks results in a characteristic GW frequency of $\,{f_{\rm GW}}\approx 1.5\,{\rm kHz}$ [right panels, see Eq. (5)].